OpenStudy (umulas):

Which of the following is a solution to the equation log 4 x + log 4 (x – 3) = 1 ?

OpenStudy (umulas):

I believe that I need to multiply it causing log4 x^2 - 3x = 1 Or is that wrong? Answers are x = 6/5 x = –3 x = 4

OpenStudy (cwrw238):

log4 ((x^2 - 3x) = 1 so (x^2 - 3x = 4^1 = 4 x^2 -3x - 4 = 0 (x-4)(x + 1) = 0 x = 4 or -1 x=4 is correct

OpenStudy (cwrw238):

log of -1 does not exist

OpenStudy (umulas):

Ok let me get this straight when we are in the part of log4 (x^2 - 3x) = 1 we multiply 4 * 1 or is it 4^1? (either way it's 1, but would want to understand) then we get x^2 - 3x = 4 x^2 - 3x - 4 (x-4)(x+1) x = 4 Correct?

OpenStudy (cwrw238):

ok the log is defined as the power that the base as to be taken to to get the number you are taking log of eg 4^2 = 16 so log4 16 = 2 4 is the base, 2 is the log

OpenStudy (cwrw238):

working the other way log4 16 = 2 4^2 = 16

OpenStudy (umulas):

Hmmmmm so it's like the power of the other side? EX: log5 125 = 5 5^5 = 125

OpenStudy (cwrw238):

thats right

OpenStudy (cwrw238):

if we have variables inlvolved log4 x = 2 then x = 4^2

OpenStudy (cwrw238):

this is 'taking logs out'

OpenStudy (umulas):

Ahhhhhh! Thanks, now I understand the question and now how to answer other questions!

OpenStudy (cwrw238):

good